# Odd dynamics of passive objects in a chiral active bath

### v1

@@ Abstract
The $\mathbb{Z}_{2}$ Steering wedge. A wedge, pictured in Fig. 1c, has no $C_{n}$ symmetry, and its $\mathbb{Z}_{2}$ reflection symmetry is broken by the bath chirality. It thus experiences net forces and torques from the bath, behaving as both a rotational and translational ratchet with average velocities $v_{\parallel}=M^{-1}\langle\bm{u}(\Theta)\cdot{\bm{P}}\rangle$ and $v_{\perp}=M^{-1}\langle\text{\bf A}\bm{u}(\Theta)\cdot{\bm{P}}\rangle$, where $\bm{u}(\Theta)=[\cos(\Theta),\sin(\Theta)]^{\mathrm{T}}$. In the achiral limit ($\ell_{g}\rightarrow\infty$) $\Omega$ and $v_{\perp}$ vanish by symmetry. In the opposite limit ($\ell_{g}\rightarrow 0$) the active force $f_{0}$ acts over a vanishing small distance, and all three ratchet currents vanish. Consequently, $\Omega$ and $v_{\perp}$ are maximal at intermediate values of $\ell_{g}$ which Fig. 4a shows to be selected for by the wedge size, $\ell_{\text{wedge}}$. Note that, while a wedge always propels towards its tip in an achiral active bath, $v_{\parallel}$ reverses at $\ell_{g}\approx\ell_{\text{wedge}}$, propelling in the opposite direction.
@@
The lack of $C_{n}$ symmetry of the wedge implies that rotational and translational motion are coupled in $\bm{\lambda}$ and $\bm{\zeta}$. This is illustrated in Fig. 4b, where the average trajectory of the wedge subject to “pulling” in the direction $\bm{u}(\Theta)$ or “pushing” in the opposite direction produces asymmetric trajectories with different directions of rotation. Unlike for the disk and rod, this coupling prevents defining effective rotational and translational temperatures for the wedge, even in the adiabatic limit, where the entropy production is positive even under the symmetrization $\omega_{0}\rightarrow-\omega_{0}$.
@@
The bath’s perspective. Decreasing the object symmetry thus drives its dynamics increasingly far from equilibrium. This raises the question as to how the broken object symmetries influence the structure and flows of the bath itself. To answer that question, we characterize the universal far-field flow generated by the object in the chiral bath, building on a procedure developed for achiral baths Baek et al. (2018); Granek et al. (2024). Far from the object, the bath dynamics is diffusive, obeying $\bm{J}_{D}=-\mathbf{D}^{\mathrm{b}}\bm{\nabla}\rho$, where $\rho(\bm{r})$ is the bath density and $\mathbf{D}^{\mathrm{b}}=D_{\parallel}^{\mathrm{b}}\bm{\delta}+D_{\perp}^{\mathrm{b}}\text{\bf A}$ is the bath diffusivity. Near the object, in contrast, the flux $\bm{J}$ of the active bath involves higher orientational moments. Defining $\bm{\delta J}=\bm{J}-\bm{J}_{D}$ and applying the steady-state continuity condition $\bm{\nabla}\cdot\bm{J}=0$ yields a Poisson equation for the density
$D_{\parallel}^{\mathrm{b}}\nabla^{2}\rho=-\bm{\nabla}\cdot\bm{\delta J}\,,$ (13)
@@
$\rho(\bm{r})=\rho_{0}+\frac{1}{2\pi\gamma D_{\parallel}}\left\{\frac{\bm{p}\cdot\bm{r}}{r^{2}}+\frac{\bm{r}\cdot\mathbf{q}\cdot\bm{r}}{2r^{4}}\right\}+\mathcal{O}(r^{-3})\,,$ (14)
@@
Here, $\bm{\nabla}V(\bm{r}_{i})=-\bm{F}_{i}$ in Eq. (1), and $\bm{m}(\bm{r})$ is the polar order field, quantifying the orientational order of the bath. Note that the integrands in Eqs. (15)-(16) vanish everywhere except at the object, where $\bm{\nabla}V\neq 0$. Plugging Eq. (14) into $\bm{J}=-\mathbf{D}^{\mathrm{b}}\bm{\nabla}\rho$ leads to an expression for the far-field flux,
$\begin{split}\bm{J}\simeq-\frac{1}{2\pi\gamma D_{\parallel}}\mathbf{D}^{\mathrm{b}}\big{(}r^{2}\bm{\delta}-2\bm{r}\otimes\bm{r}\big{)}\bigg{(}\frac{\bm{p}}{r^{4}}+\frac{2\mathbf{q}\cdot\bm{r}}{r^{6}}\bigg{)}\,,\end{split}$ (17)
where $\mathbf{q}$ is the symmetric part of $\mathbf{\tilde{q}}$. In Fig. 5, the densities and fluxes induced by the rod and wedge in both chiral and achiral active baths are succesfully compared with our multipole approximation, which show that the chirality of the bath effectively rotates the far field distribution with respect to the object. The increased departure from equilibrium as the object symmetry is decreased is reflected that the wedge is dominated by its dipole moment, while the rod’s leading contribution is only quadrupolar. Finally, the rotational flux $J_{\phi}=\bm{r}\times\bm{J}/|\bm{r}|$ of the chiral bath is observed to exhibit a large net circulation near the object. This is related to the torque on the object through Eq. (16), showing these effects to be two sides of the same coin.
Tayloring transport. Our results show how a passive object in contact with a chiral active bath turns into a chiral active particle, whose large-scale properties are determined by its shape, mass, and size. First, the object’s symmetry determines which active ratchet behaviors and transport couplings are triggered, thus determining the nature of this effective active particle. Second, we have shown how to build a Langevin description of a heavy object, that predicts effective equilibrium descriptions albeit with odd signatures of nonequilibrium. Finally, our simulations show how tuning the object size to the bath gyroradius $\ell_{g}$ maximizes the odd response coefficients and chiral ratchet motion. An exciting perspective is now to explore the interactions between multiple objects mediated by chiral active bath and the possibility of tayloring their self-assembly.

### v2

@@ Abstract
The $\Pi$ steering wedge. A wedge, pictured in Fig. 1c, has no $C_{n}$ symmetry, and the full-system parity symmetry ($\Pi$) is broken by the bath chirality. A wedge thus experiences net forces and torques from the bath, behaving as both a rotational and translational ratchet with average velocities $v_{\parallel}=M^{-1}\langle\bm{u}(\Theta)\cdot{\bm{P}}\rangle$ and $v_{\perp}=M^{-1}\langle\big{(}\text{\bf A}\bm{u}(\Theta)\big{)}\cdot{\bm{P}}\rangle$, where $\bm{u}(\Theta)=[\cos(\Theta),\sin(\Theta)]^{\mathrm{T}}$. In the achiral limit ($\ell_{g}\rightarrow\infty$) $\Omega$ and $v_{\perp}$ vanish by symmetry. In the opposite limit ($\ell_{g}\rightarrow 0$) the active force $f_{0}$ acts over a vanishing small distance, and all three ratchet currents vanish. Consequently, $\Omega$ and $v_{\perp}$ are maximal at intermediate values of $\ell_{g}$ which Fig. 4a shows to be selected for by the wedge size, $\ell_{\text{wedge}}$. Note that, while a wedge in an achiral active bath always propels towards its tip, in a chiral bath $v_{\parallel}$ reverses at $\ell_{g}\approx\ell_{\text{wedge}}$, propelling in the opposite direction.
@@
The lack of $C_{n}$ symmetry of the wedge implies that rotational and translational motion are coupled in $\bm{\lambda}$ and $\bm{\zeta}$. This is illustrated in Fig. 4b, where the average trajectory of the wedge subject to “pulling” in the direction $\bm{u}(\Theta)$ or “pushing” in the opposite direction produces asymmetric trajectories with different directions of rotation. In contrast to the disk and rod, this coupling prevents defining effective rotational and translational temperatures for the wedge, even in the adiabatic limit.
@@
The bath’s perspective. Decreasing the object symmetry thus drives its dynamics increasingly far from an effective equilibrium. This raises the question as to how the broken object symmetries influence the structure and flows of the bath itself. To answer that question, we characterize the universal far-field flow generated by the object in the chiral bath, building on a procedure developed for achiral baths Baek et al. (2018); Granek et al. (2024). Far from the object, the bath dynamics is diffusive, with current ${\bm{j}}_{D}=-\mathbf{D}^{\mathrm{b}}\bm{\nabla}\rho$, where $\rho(\bm{r})$ is the bath density and $\mathbf{D}^{\mathrm{b}}=D_{\parallel}^{\mathrm{b}}\mathbf{1}+D_{\perp}^{\mathrm{b}}\text{\bf A}$ is the bath diffusivity. Near the object, in contrast, the flux ${\bm{j}}$ of the active bath involves higher orientational moments. Defining $\bm{\delta j}={\bm{j}}-{\bm{j}}_{D}$ and applying the steady-state continuity condition $\bm{\nabla}\cdot{\bm{j}}=0$ yields a Poisson equation for the density
$D_{\parallel}^{\mathrm{b}}\nabla^{2}\rho=\bm{\nabla}\cdot\bm{\delta j}\,,$ (13)
@@
$\rho(\bm{r})=\rho_{0}+\frac{1}{2\pi\gamma D_{\parallel}}\left\{\frac{\bm{p}\cdot\bm{r}}{r^{2}}+\frac{\bm{r}\cdot\mathbf{q}\bm{r}}{2r^{4}}\right\}+\mathcal{O}(r^{-3})\,,$ (14)
@@
and $\mathbf{q}$ is the symmetric, traceless part of $\mathbf{\tilde{q}}$. Here, $\bm{F}_{i}=-\bm{\nabla}V(\bm{r}_{i})$ in Eq. (1), and $\bm{m}(\bm{r})$ is the polar order field, quantifying the orientational order of the bath. Note that the integrands in Eqs. (15)-(16) vanish everywhere except at the object, where $\bm{\nabla}V\neq 0$. Plugging Eq. (14) into ${\bm{j}}=-\mathbf{D}^{\mathrm{b}}\bm{\nabla}\rho$ for the far-field flux yields
$\begin{split}{\bm{j}}\simeq-\frac{1}{2\pi\gamma D_{\parallel}}\mathbf{D}^{\mathrm{b}}\big{(}r^{2}\mathbf{1}-2\bm{r}\bm{r}\big{)}\bigg{(}\frac{\bm{p}}{r^{4}}+\frac{2\mathbf{q}\bm{r}}{r^{6}}\bigg{)}\,.\end{split}$ (17)
In Fig. 5, the far-field densities and fluxes induced by the rod and wedge in both chiral and achiral active baths shown to agree with our multipole approximation, with chirality of the bath effectively rotating the far-field distribution with respect to the object. While the rod’s leading contribution is quadrupolar, the wedge is dominated by the dipole moment, reflecting its increased departure from equilibrium. Finally, the rotational flux $j_{\phi}=\bm{r}\times{\bm{j}}/|\bm{r}|$ of the chiral bath is observed to exhibit a large net circulation near the object. This is related to the torque on the object through Eq. (16), showing these effects to be two sides of the same coin.
Tailoring transport. Our results show how a passive object in contact with a chiral active bath is imbued with odd transport and ratchet properties determined by its shape, mass, and size. First, the object’s symmetry determines which active ratchet behaviors and transport couplings are triggered. Second, we have shown how to build a Langevin description of a heavy object, predicting effective equilibrium descriptions that nonetheless retain odd signatures of nonequilibrium. Finally, our simulations show how tuning the object size to the bath gyroradius $\ell_{g}$ maximizes the odd response coefficients and chiral ratchet motion. An exciting prospect is now to explore the interactions between multiple objects mediated by chiral active bath and the possibility of tailoring their self-assembly.